OpenStudy (nptweezy):

ASAP Please I'm trying to find out if x^4 is the reciprocal of x^-4 or is it vice versa

OpenStudy (mathstudent55):

What is the product of reciprocals?

OpenStudy (mathstudent55):

Ok, let's go back to the definition. Reciprocals are number whose product is 1.

OpenStudy (mathstudent55):

If you have two numbers that when you multiply them together you get 1, then the numbers are reciprocals.

OpenStudy (mathstudent55):

For example: Question: are 2 and 1/2 reciprocals? Answer: Just multiply them together and find the result. If the result is 1, then 2 and 1/2 are reciprocals. 2 * 1/2 = 2/1 * 1/2 = 2/2 = 1/1 = 1 Since the product of 2 and 1/2 is 1, 2 and 1/2 are reciprocals.

OpenStudy (nptweezy):

Thanks I get now

OpenStudy (mathstudent55):

Now multiply the two expressions you are given. If the result is 1, then they are reciprocals.

OpenStudy (mathstudent55):

Wlcm